OpenStudy (anonymous):

A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. The probability that you have 1 cherry candy and 2 lemon candies is

OpenStudy (anonymous):

\[ \huge \frac{\binom 6 1 \times \binom 2 2 }{\binom {11} 3} \]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\frac 2 {55} \]